10. A physics student is trying to impress people by working out at the gym. a) How much Gravitational Potential Energy would the student impart to a 35.0 kg weight in raising it from the floor tol.50 m high? 12 marks] b) How much mechanical work is done by the student on the weight? 12 marks] What is the power output of the student if raising the weight took 1.23 s? 12 marks] c) d) How much mechanical work is done by the student on the weight when the student lowers the weight from 1.50 m down to the floor? [2 marks
Answer #1

here,

mass , m = 35 kg

height , h0 = 1.5 m

10)

the gravitational potential energy , PEg = m * g * h

PEg = 35 * 9.81 * 1.5 J

PEg = 515.025 J

b)

the mechanical work done by the student on weight , W = PEg = 515.025 J

c)

time taken , t = 1.23 s

teh power output , P = W/t

P = 515.025 /1.23 W = 418.7 W

d)

the work done by the student , W' = m * g * h

W' = 35 * 9.81 * 1.5 = 515.025 J